A particle has shifted along some trajectory in the plane xy from point 1 whose radius vector r1 = i + 2j to point 2 with the radius vector r2 = 2i - 3j. During that time the particle experienced the action of certain forces, one of which being F = 3i + 4j. Find the work performed by the force F. (Here r1, r2, and F are given in SI units).
1
Expert's answer
2020-10-15T10:41:25-0400

"A=|\vec F|\cdot |\vec s|\cdot cos(φ)=\vec F\cdot \vec s" ,

where "|\vec F|" (N, Newton) is the vector "\vec F" length, "|\vec s|" (m, meter) is the vector "\vec s" length

( "\vec s=\vec r_{2}-\vec r_{1}=(2\vec i-3\vec j)-(\vec i+2\vec j)=\vec i-5\vec j" ), "\u03c6" is the angle between "\vec F" and "\vec s" .

Thereby "|\vec F|\cdot cos(φ)" is the length of the projection of the vector "\vec F" to the vector "\vec s" .

"A=\vec F\cdot \vec s=(3\vec i+4\vec j)(\vec i-5\vec j)="

"=3\vec i^2-15\vec i\vec j+4\vec j\vec i-20\vec j^2=3-0+0-20="

"=3-20=-17J" (Joule).

Answer: −17J.